9. O-13 points CJ9 18.P.O54 A spherical surface completely surrounds a collection of charges. Find the electric flux through the surface if the collection consists of the following charges. (a) a single +3.65 x 10-6 C charge N. m2/C (b) a single-2.40 x 10-6 C charge N m2/c (c) both of the charges in (a) and (b) N m2/c 10. O-1 points C.J9 18 P057 sc 名0 F6 19
Answer #1

a) electric flux = Q / e0

= 3.65 * 10-6 / (8.85 * 10-12)

= 4.12 * 105 N.m2/C

b) electric flux = 2.4 * 10-6 / (8.85 * 10-12)

= 2.71 * 105 N.m2/C

c) electric flux = ((3.65-2.4) * 10-6) / (8.85 * 10-12)

= 1.41 * 105 N.m2/C